OpenStudy (anonymous):

Finding the binomial probabilities of: n=11. p=.5, P(x>8) So do I find P(x=0)+P(x=1)...+P(x=7) in the binomial probability formula?

OpenStudy (aum):

But that would be finding P(x < 8). They want P(x > 8).

OpenStudy (anonymous):

So do I find 8 -11?

OpenStudy (aum):

P(x=9) + P(x=10) + P(x=11)

OpenStudy (anonymous):

little confusing but thank you

OpenStudy (aum):

If you prefer you can find P(x=0)+P(x=1)...+P(x=7) and then calculate 1 - the sum of probabilities.

OpenStudy (anonymous):

Thank you for the help...I had P(x=0)+...P(x=7) all figured but i erased, and i wished i would have thought that.

OpenStudy (loser66):

what does p =0.5 mean?

OpenStudy (anonymous):

P is probability of success or failures in the binomial

OpenStudy (loser66):

and what is the probability of mass function? f(x) =?

OpenStudy (loser66):

oh, you mean the probability of failure or success is 1/2.

OpenStudy (anonymous):

yep

OpenStudy (aum):

With P(x=0)+P(x=1)...+P(x=7) you have to calculate 8 probabilities and add then up and then subtract the sum from 1. But the easier method would be P(x=9) + P(x=10) + P(x=11) where you calculate 3 probabilities and add them up.